The following two tables represent the price and quantity of the goods produced within an economy in 2015 & 2016.

What is the inflation rate between 2015 & 2016?

Prices Quantities
2015 2016 2015 2016
Soda $2 $2 36 40
Burgers $8 $12 8 10
French Fries $2 $1 16 20
Answer #1

Find nominal GDP in 2015 (which is also the real GDP of 2015 since 2015 is the base year) = price of soda (2015) x quantity of soda (2015) + price of burgers (2015) x quantity of burgers (2015) + price of french fries (2015) x quantity of french fries (2015) = 2*36 + 8*8 + 2*16 = 168

Find nominal GDP in 2016 = price of soda (2016) x quantity of soda (2016) + price of burgers (2016) x quantity of burgers (2016) + price of french fries (2016) x quantity of french fries (2016) = 2*40 + 12*10 + 1*20 = 220

Find real GDP in 2016 = price of soda (2015) x quantity of soda (2016) + price of burgers (2015) x quantity of burgers (2016) + price of french fries (2015) x quantity of french fries (2016) = 2*40 + 8*10 + 2*20 = 200

Now GDP deflator 2015 = nominal GDP 2015 / Real GDP 2015 = 168/168 = 100

GDP deflator 2016 = nominal GDP 2016 / Real GDP 2016 = 220/200 = 110

Inflation rate = (110 - 100)/100 = 10%

Therefore the rate of inflation is 10%